OpenStudy (anonymous):

I need to find the center,radius, and then graph this x^2+y^2+10y+21=0

OpenStudy (anonymous):

I know the formula would be (x-a)^2(y-b)^2=r^2

OpenStudy (shubhamsrg):

Whatever you said above was correct. (h,k) is the center of (x-h)^2 + (y-k)^2 = r^2 now when you expand all the terms, x^2 -2xh + h^2 + y^2 - 2yk + k^2 = r^2 x^2 + y^2 -2hx -2ky + h^2 + k^2 - r^2 =0 Your given equation is of this form only. Now recall that (h,k) was the center that means if the expression is represented like this, then negative halves of the coefficients of x and y are the respective co-ordinates of the center of the circle. for example, in x^2 + y^2 - 2x + 4y + 2 = 0 (-(-2)/2 , -(4)/2 ) are the co-ordinates of the center i.e. (1,-2) is the center now, using this, can you tell the center of your given equation ?

OpenStudy (loser66):

@shubhamsrg gave you a complete explanation, you still pump it. where are you stuck?

OpenStudy (loser66):

make questions, one by one, please. If I can, I will explain. If I cannot, I will tag other for help.

OpenStudy (anonymous):

Ok I am now stuck at x^2 + y^2 -2hx -2ky + h^2 + k^2 - r^2 =0

OpenStudy (loser66):

you don't know why you need that form, right?

OpenStudy (anonymous):

to find the center and radius?

OpenStudy (loser66):

yes.

OpenStudy (loser66):

because that form is expanding form of (x-h)^2 +(y-k)^2 = something.

OpenStudy (loser66):

ok, break it down. expand this first: (x-h)^2 = ?

OpenStudy (anonymous):

x^2-h^2

OpenStudy (loser66):

nope, friend, not that. again, try !! if you don't know, google. by that way, you remember your mistake.

OpenStudy (loser66):

I can give you, Sure!! but I want you find out by yourself

OpenStudy (anonymous):

what do you mean expand?

OpenStudy (anonymous):

ooooooooo

OpenStudy (anonymous):

wait I got it

OpenStudy (anonymous):

(x-h)(x-h)

OpenStudy (loser66):

(x -h)^2 = x^2 -2xh +h^2 . That's called "expand"

OpenStudy (loser66):

yes!! your way is ok, just one more step to get mine

OpenStudy (anonymous):

oh well I was off

OpenStudy (loser66):

and now , expand (y-k)^2 = ?

OpenStudy (anonymous):

(y-k)^2=y^2-2yk+h^2

OpenStudy (loser66):

k^2 not h^2, right?

OpenStudy (anonymous):

(y-k)^2=y^2-2yk+k^2 right

OpenStudy (loser66):

now combine them, (x-h)^2 +(y-k)^2 =? (expand form)

OpenStudy (anonymous):

(x -h)^2 = x^2 -2xh +h^2+(y-k)^2=y^2-2yk+k^2

OpenStudy (loser66):

yeah!!! now compare it with yours, what they are.

OpenStudy (anonymous):

x^2 -2xh +h^2+y^2-2yk+k^2

OpenStudy (anonymous):

sorta not really, like I see that they have some similarities

OpenStudy (loser66):

ok